What minimum speed does a 190 g puck need to make
Answer #1

given data is m = 190 g = 0.190 kg

length L = 3.2 m

angle \theta = 290

we have h = L sin\theta

here h is height of th ramp

h = 3.2 X sin29

h = 1.55 m

and conservation of energy that

or equilibrium is KE = PE

1 / 2 m v2 = m g h

h = v2 / 2 g

1.55 = v2 / 2 X 9.8

v2 = 2 X 9.8 X 1.55

v2 = 30.38

v = 5.511 m/s